OpenStudy (anonymous):

solve integration of e^-3x from 0 to 5?

OpenStudy (amistre64):

youre missing a -3

OpenStudy (anonymous):

nothing is missing :)

OpenStudy (amistre64):

..... youre missing a -3; so borrow it from -3/-3

OpenStudy (amistre64):

you know how to integrate -3 e^-3x i assume

OpenStudy (amistre64):

\[\int e^{-3x}dx=\ -\frac{1}{3}\int -3e^{-3x}dx\]

OpenStudy (anonymous):

Good ol' Integration. :)

OpenStudy (anonymous):

@amistre yea yea I got it.. thanks ;) can I reward you?

OpenStudy (lalaly):

give him a medal hit a gd answer

OpenStudy (amistre64):

you can reward me if you want, but a simple thanks is really all i require :)